Grade 12th passGeneral Physics

in the double slit experiment if one slit have a width as twice as the other what will happens to the interference pattern

moon watcher
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago

Maximum amplitude & Intensity -

When \theta = 0,2\pi ---2n\pi
 

- wherein

A_{max }= A_{1}+A_{2}

I_{max }= \left ( \sqrt{I_{1}}+\sqrt{I_{2}} \right )^{2}

 

 and

 

Minimum amplitude & Intensity -

\theta = \left ( 2n+1 \right )\pi
 

- wherein

A_{min }= A_{1}-A_{2}

I_{min }= \left ( \sqrt{I_{1}}-\sqrt{I_{2}} \right )^{2}

 

 In this case, since A \propto \ width of slits 

A_{1} = 2A

A_{2} = A

A^{max} = A_{1}+A_{2} = 3A and  A^{min} = \left ( A_{1} \right-A_{2} ) =A

Hence amplitude/intensity of both maxima and minima increases